An 8-bit register contains the Hex value 58. What is the register value after an arithmetic shift right? Starting from the initial number 58, determine the register value i) after logical shift right ii) after an arithmetic shift left, and state whether there is an overflow not, explain through the proper diagrams.
1
Expert's answer
2021-09-26T11:56:48-0400

V = 3,510,593 = 0x 00359141 V = 0000,0000,0011,0101,1001,0001,0100,0001 = 1.10101100100010100000 ×2^(21) = (-1)s × M × 2^E s = 0; E =21=exp - Bias => exp = E+Bias = 21+127=148 = 128+16+4 exp = 1001,0100 M = 1+ frac×2^(-23) = 1.f22f21¼f1f0 = 1.10101100100010100000 Hence frac = 101,0110,0100,0101,0000,0100

 

 Its IEEE single-precision representation is s exp frac 0 1001,0100 101,0110,0100,0101,0000,0100 which is 0, 1001,0100, 101,0110,0100,0101,0000,0100 which is regrouped in 4 bits i.e., 0100,1010,0101,0110,0100,0101,0000,0100 in hexadecimal format is 0x4A564504